What is the kinetic energy of a


proton (proper mass is m0 = 1.67 x 10^-27 kg) travelling at 40 % of the speed


of light?

1
Expert's answer
2022-02-04T18:00:30-0500

"\gamma=\frac{1}{\sqrt{1-(v/c)^2}}=\frac{1}{\sqrt{1-(0.4c/c)^2}}=\frac{1}{\sqrt{1-0.16}}=1.091"


"E_k=(\gamma-1)mc^2=(1.091-1)\times1.67\times10^{-27}kg\times(2.998\times10^8\frac{m}{s})^2=1.37\times10^{-11}J"


Answer: "1.37\times10^{-11}J"